# The Wavefront Set: Bounds for the Langlands parameter

### v2

@@ 3.1. Regular supercuspidal representations
For all regular depth-zero supercuspidal representations [KV, DBR], the conjecture is trivially true since in that case ${\mathfrak{h}}^{\vee}$ is just a Cartan subalgebra of $\mathfrak{g}^{\vee}$, so the only possible Langlands nilpotent element is 00. Hence the conjecture just says that ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq{{\mathcal{O}}}_{p}$, where ${{\mathcal{O}}}_{p}$ is the principal ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit, which is true, but trivial. On the other hand, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={{\mathcal{O}}}_{p}$ if and only if $\pi$ is a generic representation, and this is not always the case. In general, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ can be equal to the ${\mathbf{G}}(\bar{\mathsf{k}})$-saturation of the principal nilpotent orbit in a maximal pseudo-Levi, as we explain next.
@@
The element $w^{\vee}\theta^{\vee}$ is the image of the inverse Frobenius under the Langlands homomorphism $\varphi$. Then a complete (enhanced) Langlands parameter is given by a triple $(\lambda,w,\rho)$, where $\rho$ is an irreducible representation of the finite group ${T^{\vee}}^{w^{\vee}\theta^{\vee}}=Z_{G^{\vee}}(\varphi).$ If $X$ is the weight lattice of ${\mathbf{G}}$, and $W$ is the Weyl group of ${\mathbf{G}}$, $\theta$ the Galois automorphism giving $\theta^{\vee}$, there is a natural isomorphism
$X/(1-w\theta)X\longleftrightarrow\mathsf{Irr}~{}{T^{\vee}}^{w^{\vee}\theta^{\vee}},\quad x\mapsto\rho_{x}.$ (3.3)
By abuse of notation, denote $x$ also the preimage in $X$ of an element in $X/(1-w\theta)X$. Let $\pi(\lambda,w,0,x)$ denote the supercuspidal representation attached to $(\lambda,w,x)$ in loc. cit. The 00 here stands for the zero nilpotent orbit, i.e., the image of the $SL(2)$ under the Langlands parameter must be trivial (since ${\mathfrak{g}^{\vee}}^{\lambda(I_{\mathsf{k}})}=\mathfrak{t}^{\vee}$).
@@
$\pi(\lambda,w,0,x)=\operatorname{ind}_{Z^{F_{x}}{\mathbf{G}}_{x}^{F_{x}}}^{{\mathbf{G}}^{\omega_{x}}({\mathsf{k}})}(\kappa_{x}).$
@@
4. Depth-zero representations of classical groups
In [LuSt], Lust and Stevens constructed L-packets of depth-zero supercuspidal representations of classical groups. We will verify Conjecture 1.6 for these $L$-packets. Let ${\mathcal{G}}$ be a connected reductive group of classical type with connected center over ${\mathbb{F}}_{q}$, $q$ odd. Let $F$ be the Frobenius map. Let ${\mathcal{G}}^{*}$ be the dual group over ${\mathbb{F}}_{q}$ with dual Frobenius also denoted $F$. Let $\mathsf{Irr}{\mathcal{G}}^{F}$ be the set of equivalence classes of irreducible complex ${\mathcal{G}}^{F}$-representations. According to Lusztig’s classification, we can partition
@@ 4. Depth-zero representations of classical groups
where ${\mathcal{G}}_{s}$ is the endoscopic group of ${\mathcal{G}}$ with dual ${\mathcal{G}}^{*}_{s}:=Z_{{\mathcal{G}}^{*}}(s)$. To have cuspidal representations, we need $Z({\mathcal{G}}^{*})^{0}$ and $Z({\mathcal{G}}_{s}^{*})^{0}$ to have the same $F$-rank. In this case, $\Psi_{s}^{{\mathcal{G}}}$ gives a correspondence of cuspidal represenattions. In other words, the set of cuspidal ${\mathcal{G}}^{F}$-representations are in one-to-one correspondence with ${\mathcal{G}}^{*}$-pairs $(s,\tau)$ where $s\in{{\mathcal{G}}^{*}}^{F}$ is semisimple of equal rank and $\tau$ is a cuspidal unipotent representation of ${\mathcal{G}}_{s}^{F}$.
@@
    if $P(x)\neq x\pm 1$, then $a_{P}=\frac{1}{2}(m_{P}^{2}+m_{P})$ for some integer $m_{P}\geq 0.$
@@
    if we set $a_{\pm}=a_{x\mp 1}$, there are some conditions on $a_{\pm}$ depending on the type of the group, see [LuSt, §7.2].
@@
Suppose ${\mathcal{G}}=Sp(2n)$. If $P(x)$ appears in $P_{s}(x)$ and $\deg(P)>1$, then necessarily $\deg(P)$ is even. We denote $\deg(P)=2k_{P}$, $k_{P}>0$ integer. Otherwise,
$a_{+}=2(m_{+}^{2}+m_{+})+1,\quad a_{-}=2m_{-}^{2},$
for some nonnegative integers $m_{\pm}$. Then a cuspidal representation of $Sp(2n,{\mathbb{F}}_{q})$ is uniquely determined by a collection of nonnegative integers $m_{\pm}$, $m_{P},k_{P}$ such that
@@ 4.1. Nilpotent Langlands parameters
Lust and Stevens [LuSt, §8] attach to $\pi$ the nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ given by the partition obtained by the following procedure. For every integer $\alpha$ denote by $[\alpha]$ the string $\{\alpha,\alpha-2,\alpha-4,\dots,1\}$ if $\alpha>0$ is odd, $\{\alpha,\alpha-2,\alpha-4,\dots,2\}$ if $\alpha>0$ is even, and $[\alpha]=\emptyset$ if $\alpha\leq 0$. Start with an empty partition and
@@
    if $k_{P}>0$, append $2k_{P}$ copies of $[m_{P}^{(1)}+m_{P}^{(2)}]$ and $2k_{P}$ copies of $[|m_{P}^{(1)}-m_{P}^{(2)}|-1]$.
@@
    append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})+1]$ and $[2|m_{+}^{(1)}-m_{+}^{(2)}|-1]$ if $n_{2}^{an}=1$, or the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})]$ and $[2|m_{+}^{(1)}-m_{+}^{(2)}+1|-1]$ if $n_{2}^{an}$ is even;
@@
    append the strings $[2(m_{-}^{(1)}+m_{-}^{(2)})]$ and $[2|m_{-}^{(1)}-m_{-}^{(2)}-1|-1]$ if $n_{2}^{an}=1$, or the strings $[2(m_{-}^{(1)}+m_{-}^{(2)})-1]$ and $[2|m_{-}^{(1)}-m_{-}^{(2)}|-1]$ if $n_{2}^{an}$ is even.
@@
    • If $n_{1}^{an}=1$ and $n_{2}^{an}$ is even, append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})]$ and $[|2(m_{+}^{(1)}-m_{+}^{(2)})+1|-1]$.
    • If $n_{1}^{an}$ and $n_{2}^{an}$ are both even, append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})-1]$ and $[|2(m_{+}^{(1)}-m_{+}^{(2)})|-1]$.
    • If $n_{1}^{an}=n_{2}^{an}=1$, append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})+1]$ and $[|2(m_{+}^{(1)}-m_{+}^{(2)})|-1]$.
@@ 4.2. Symplectic groups
Definition 4.2.
If $\alpha,\beta$ are two partitions of $m,n$, respectively, the sum $\alpha+\beta$ is a partition of $m+n$ given as follows. Suppose $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{\ell})$ and $\beta=(\beta_{1},\beta_{2},\dots,\beta_{\ell})$ with $0\leq\alpha_{1}\leq\alpha_{2}\leq\dots\leq\alpha_{\ell}$, $0\leq\beta_{1}\leq\beta_{2}\leq\dots\leq\beta_{\ell}$. (By allowing entries of 00, we may assume there is the same number of parts in $\alpha$ and $\beta$.) Then $\alpha+\beta=(\alpha_{1}+\beta_{1},\alpha_{2}+\beta_{2},\dots,\alpha_{\ell}+\beta_{\ell}).$
This operation on partitions plays a key role in the explicit computation of $j$-induction for representations of the Weyl group of classical types.
At one extreme suppose that only endoscopic unitary factors appear, that is only polynomials $P$ with $\deg P>1$. Hence ${\mathbb{O}}^{\vee}_{\pi}$ is obtained by appending $2k_{P}$ copies of $[m_{P}^{(1)}+m_{P}^{(2)}]$ and $2k_{P}$ copies of $[|m_{P}^{(1)}-m_{P}^{(2)}|-1]$, for each $k_{P}>0$. On each $Sp(2n_{i})$, $i=1,2$, factor, we have
@@
$j_{\prod_{P}S_{a_{P}}^{k_{P}}}^{W_{n}}\left(\boxtimes_{P}(1,2,\dots,m_{P})^{\boxtimes k_{P}}\right).$
@@
$j_{\prod_{P}S_{k_{P}a_{P}}}^{W_{n}}\left(\boxtimes_{P}(k_{P},2k_{P},\dots,m_{P}k_{P})\right)=j_{\prod_{P}W_{k_{P}a_{P}}}^{W_{n}}(\sigma_{P}),$ (4.3)
@@
$\sigma_{P}=\begin{cases}(k_{P},3k_{P},\dots,m_{P}k_{P})\times(2k_{P},4k_{P},\dots,(m_{P}-1)k_{P}),&m_{P}\text{ odd}\\ (2k_{P},4k_{P},\dots,m_{P}k_{P})\times(k_{P},3k_{P},\dots,(m_{P}-1)k_{P}),&m_{P}\text{ even}.\end{cases}$
@@
$j_{\prod_{P}S_{a_{P}}^{k_{P}}}^{W_{n}}\left(\boxtimes_{P}(1,2,\dots,m_{P})^{\boxtimes k_{P}}\right)$
@@
By the calculation leading up to (4.4), we need to show that the nilpotent orbit for $j_{\prod_{P}W_{k_{P}a_{P}}}^{W_{n}}(\boxtimes_{P}\sigma_{P})$ is as claimed in the statement of the lemma. Since $j$-induction works in stages, this comes down to the case of two polynomials $P_{1},P_{2}$, that is to the easy verification that $j_{W_{1}\times W_{2}}^{W}(\sigma_{1}\boxtimes\sigma_{2})$ is the Springer representation for the nilpotent $(2k_{1},4k_{1},\dots,2m_{1}k_{1})+(2k_{2},4k_{2},\dots,2m_{2}k_{2})$. ∎
@@
    ${\mathbb{O}}^{\vee}_{\pi}=\bigsqcup_{P}[m_{P}^{(1)}+m_{P}^{(2)}]^{2k_{P}}\sqcup[|m_{P}^{(1)}-m_{P}^{(2)}|-1]^{2k_{P}}$.
@@
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\bigsqcup_{i=1,2}\left(\sum_{P}(2k_{P},4k_{P},\dots,2m_{P}^{(i)}k_{p})\right)$
@@
Claim (1) is just a restatement of the Lust-Stevens parametrization. Claim (2) is directly implied by Lemma 4.3. It remains to verify (3).
@@
${\mathbb{O}}^{\vee}_{\pi}=\bigsqcup_{P}(\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}}),$
@@
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\sum_{P}(2k_{P},4k_{P},\dots,2m_{P}k_{p})$
Notice that
$d^{\vee}((\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}}))$ $=(\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}})^{t}$
$=(2k_{P},4k_{P},\dots,2m_{P}k_{p})$
which proves that $d^{\vee}({\mathbb{O}}_{\pi}^{\vee})={}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ if there is a single $P$. If there is more than one polynomial $P$, the claim follows as well by the compatibility
$(\alpha\sqcup\beta)^{t}=\alpha^{t}+\beta^{t}$ (4.5)
for two (or more) partitions $\alpha,\beta$.
@@
$d^{\vee}([m^{(1)}+m^{(2)}]^{2k}\sqcup[|m^{(1)}-m^{(2)}|-1]^{2k})$ $=\left((m^{(1)}+m^{(2)},m^{(1)}+m^{(2)}-2,\dots,m^{(1)}-m^{(2)}+2\right.,$
@@
$=(2k,4k,\dots,2m^{(1)}k)\sqcup(2k,4k,\dots,2m^{(2)}k),$
@@
$=\sum_{P}\left((2k_{P},4k_{P},\dots,2m_{P}^{(1)}k_{P})\sqcup(2k_{P},4k_{P},\dots,2m_{P}^{(2)}k_{P})\right)$
It is easy to see that this is greater in the partition order than ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\bigsqcup_{i=1,2}\left(\sum_{P}(2k_{P},4k_{P},\dots,2m_{P}^{(i)}k_{p})\right)$. Notice that the difference between the two formulae is simply that $\sigma$ and $\sqcup_{i=1,2}$ are interchanged.

### v3

@@ 3.1. Regular supercuspidal representations
For all regular depth-zero supercuspidal representations [KV, DBR], the conjecture is trivially true since in that case $\mathfrak{h}$$^{\vee}$ is just a Cartan subalgebra of $\mathfrak{g}^{\vee}$, so the only possible Langlands nilpotent element is 0. Hence the conjecture just says that $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq{{\mathcal{O}}}_{p}$, where ${{\mathcal{O}}}_{p}$ is the principal ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit, which is true, but trivial. On the other hand, $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={{\mathcal{O}}}_{p}$ if and only if $\pi$ is a generic representation, and this is not always the case. In general, $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ can be equal to the ${\mathbf{G}}(\bar{\mathsf{k}})$-saturation of the principal nilpotent orbit in a maximal pseudo-Levi, as we explain next.
@@
The element $w^{\vee}\theta^{\vee}$ is the image of the inverse Frobenius under the Langlands homomorphism $\varphi$. Then a complete (enhanced) Langlands parameter is given by a triple $(\lambda,w,\rho)$, where $\rho$ is an irreducible representation of the finite group $(T^{\vee})^{w^{\vee}\theta^{\vee}}=Z_{G^{\vee}}(\varphi).$ If $X$ is the weight lattice of $\mathbf{G}$}, and $W$ is the Weyl group of $\mathbf{G}$}, $\theta$ the Galois automorphism giving $\theta^{\vee}$, there is a natural isomorphism
$X/(1-w\theta)X\longleftrightarrow\mathsf{Irr}~(T^{\vee})^{w^{\vee}\theta^{\vee}},\quad x\mapsto\rho_{x}.$ (3.3)
By abuse of notation, denote $x$ also the preimage in $X$ of an element in $X/(1-w\theta)X$. Let $\pi(\lambda,w,0,x)$ denote the supercuspidal representation attached to $(\lambda,w,x)$ in loc. cit. The 0 here stands for the zero nilpotent orbit, i.e., the image of the $SL(2)$ under the Langlands parameter must be trivial (since $(\mathfrak{g}^{\vee})^{\lambda(I_{\mathsf{k}})}=\mathfrak{t}^{\vee}$).
@@
$\pi(\lambda,w,0,x)=\operatorname{ind}_{\mathbf{Z}^{F_{x}}{\mathbf{G}}_{x}^{F_{x}}}^{{\mathbf{G}}^{\omega_{x}}({\mathsf{k}})}(\widetilde{\kappa}_{x}).$
@@
4. Depth-zero supercuspidal representations of classical groups
In [LuSt], Lust and Stevens constructed $L$-packets of depth-zero supercuspidal representations of classical groups. We will verify Conjecture 1.6 for these $L$-packets. Let $\mathcal{G}$} be a connected reductive group of classical type over ${\mathbb{F}}_{q}$, $q$ odd. Let $F$ be the Frobenius map. Let $\mathcal{G}$$^{*}$ be the dual group over ${\mathbb{F}}_{q}$ with dual Frobenius also denoted $F$. Let $\mathsf{Irr}\,{\mathcal{G}}^{F}$ be the set of equivalence classes of irreducible complex ${\mathcal{G}}^{F}$-representations. According to Lusztig’s classification, we can partition
@@ 4. Depth-zero supercuspidal representations of classical groups
[…]
@@
    if $P(x)\neq x\pm 1$, then it has even degree $n_{p}=2k_{P}$, $k_{P}\geq 1$, and multiplicity $a_{P}=\frac{1}{2}(m_{P}^{2}+m_{P})$ for some integer $m_{P}\geq 0.$
@@
    if we set $a_{\pm}=a_{x\mp 1}$, there are integers $m_{\pm}\geq 0$ with conditions, depending on the type of the group, see [LuSt, §7.2], as follows:
    1. (a)
      if ${\mathcal{G}}=SO(2n+1)$, then $a_{+}=2(m_{+}^{2}+m_{+})$ and $a_{-}=2(m_{-}^{2}+m_{-})$;
    2. (b)
      if ${\mathcal{G}}=GSO^{\pm}(2n)$, then $a_{+}=2m_{+}^{2}$ and $a_{-}=2m_{-}^{2}$;
    3. (c)
      if ${\mathcal{G}}=Sp(2n)$, then $a_{+}=2(m_{+}^{2}+m_{+})+1$ and $a_{-}=2m_{-}^{2}$.
@@
A cuspidal representation of $Sp(2n,{\mathbb{F}}_{q})$ is uniquely determined by a collection of nonnegative integers $m_{\pm}$, $m_{P},k_{P}$ such that
@@ 4.1. Nilpotent Langlands parameters
Lust and Stevens [LuSt, §1 and §8] attach to $\pi$ the nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ given by the partition obtained by the following procedure. For every integer $\alpha$ denote by $\alpha]$ the string $\{\alpha,\alpha-2,\alpha-4,\dots,1\}$ if $\alpha>0$ is odd, $\{\alpha,\alpha-2,\alpha-4,\dots,2\}$ if $\alpha>0$ is even, and $\alpha]=\emptyset$ if $\alpha\leq 0$. Start with an empty string and repeat the following procedure for each $P$:
@@
    if $n_{P}\geq 2$, append $n_{P}$ copies of $m_{P}^{(1)}+m_{P}^{(2)}]$ and $n_{P}$ copies of $|m_{P}^{(1)}-m_{P}^{(2)}|-1]$.
@@
    append the strings $2(m_{+}^{(1)}+m_{+}^{(2)})+1]$ and $2|m_{+}^{(1)}-m_{+}^{(2)}|-1]$ if $n_{2}^{an}=1$, or the strings $2(m_{+}^{(1)}+m_{+}^{(2)})]$ and $2|m_{+}^{(1)}-m_{+}^{(2)}+1|-1]$ if $n_{2}^{an}$ is even;
@@
    append the strings $2(m_{-}^{(1)}+m_{-}^{(2)})]$ and $2|m_{-}^{(1)}-m_{-}^{(2)}-1|-1]$ if $n_{2}^{an}=1$, or the strings $2(m_{-}^{(1)}+m_{-}^{(2)})-1]$ and $2|m_{-}^{(1)}-m_{-}^{(2)}|-1]$ if $n_{2}^{an}$ is even.
@@
    • If $n_{1}^{an}=1$ and $n_{2}^{an}$ is even, append the strings $2(m_{+}^{(1)}+m_{+}^{(2)})]$ and $|2(m_{+}^{(1)}-m_{+}^{(2)})+1|-1]$.
    • If $n_{1}^{an}$ and $n_{2}^{an}$ are both even, append the strings $2(m_{+}^{(1)}+m_{+}^{(2)})-1]$ and $|2(m_{+}^{(1)}-m_{+}^{(2)})|-1]$.
    • If $n_{1}^{an}=n_{2}^{an}=1$, append the strings $2(m_{+}^{(1)}+m_{+}^{(2)})+1]$ and $|2(m_{+}^{(1)}-m_{+}^{(2)})|-1]$.
@@
4.2. Partitions
We record several elementary facts about partitions that will be used in the proofs in the next subsections.
Definition 4.2.
If $\alpha,\beta$ are two partitions of $m,n$, respectively, the sum $\alpha+\beta$ is a partition of $m+n$ given as follows. Suppose $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{\ell})$ and $\beta=(\beta_{1},\beta_{2},\dots,\beta_{\ell})$ with $\alpha_{1}\geq\alpha_{2}\geq\dots\geq\alpha_{\ell}\geq 0$, $\beta_{1}\geq\beta_{2}\geq\dots\geq\beta_{\ell}\geq 0$. (By allowing entries of 0, we may assume there is the same number of parts in $\alpha$ and $\beta$.) Then $\alpha+\beta=(\alpha_{1}+\beta_{1},\alpha_{2}+\beta_{2},\dots,\alpha_{\ell}+\beta_{\ell}).$
The above operation on partitions plays a key role in the explicit computation of $j$-induction for representations of the Weyl group of classical types.
If $\alpha$ and $\beta$ are two partitions, denote by $\alpha\sqcup\beta$ the partition obtained by concatenation (and reordering). If $\alpha$ is a partition, let $\alpha^{t}$ denote the transpose partition. Let $\alpha\leq\beta$ denote the dominance order of partitions.
Lemma 4.3.
  1. (1)
    If $\alpha\leq\beta$ then $\alpha^{t}\geq\beta^{t}$.
  2. (2)
    If $\alpha_{1}\leq\beta_{1}$ and $\alpha_{2}\leq\beta_{2}$ then $\alpha_{1}\sqcup\alpha_{2}\leq\beta_{1}\sqcup\beta_{2}$.
  3. (3)
    For any two partitions $\alpha,\beta$,
    $(\alpha\sqcup\beta)^{t}=\alpha^{t}+\beta^{t}$ (4.3)
  4. (4)
    For any four partitions $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$,
    $(\alpha_{1}\sqcup\beta_{1})+(\alpha_{2}\sqcup\beta_{2})\geq(\alpha_{1}+\alpha_{2})\sqcup(\beta_{1}+\beta_{2}).$ (4.4)
  5. (5)
    For any three partitions $\alpha,\beta,\gamma$, if $\alpha\geq\beta$, then $\alpha+\gamma\geq\beta+\gamma$.
Proof.
Claims (1), (2), (3) are elementary and well known, see [CMBO2, §5.1] for example. Claim (5) follows from (2) and (3). We prove (4).
Let us first prove the case of three partitions:
$\beta_{1}+(\alpha\sqcup\beta_{2})\geq\alpha\sqcup(\beta_{1}+\beta_{2}).$
by induction on the length of $\alpha$. If $\alpha=\emptyset$, this is clearly true. Suppose $\alpha=\{s\}$ is a singleton. Write $\beta_{1}=(a_{1}\geq a_{2}\geq\dots\geq a_{k})$ and $\beta_{2}=(b_{1}\geq b_{2}\geq\dots\geq b_{k})$ and suppose $i$ and $t$ are the smallest indices such that $b_{i}>s\geq b_{i+1}$ and $a_{t}+b_{t}>s\geq a_{t+1}+b_{t+1}$. (We think $a_{0}=b_{0}=\infty$ for the consistency of the notation.) Then
$\beta_{1}+(\beta_{2}\sqcup\{s\}):\ a_{1}+b_{1}\geq a_{2}+b_{2}\geq\dots\geq a_{i}+b_{i}\geq a_{i+1}+s\geq a_{i+2}+b_{i+1}\geq\dots\geq a_{k}+b_{k-1}\geq b_{k},$
while
$(\beta_{1}+\beta_{2})\sqcup\{s\}:\ a_{1}+b_{1}\geq a_{2}+b_{2}\geq\dots\geq a_{t}+b_{t}>s\geq a_{t+1}+b_{t+1}\geq\dots\geq a_{k}+b_{k}.$
and it is immediate that $\beta_{1}+(\beta_{2}\sqcup\{s\})\geq(\beta_{1}+\beta_{2})\sqcup\{s\}$.
Now suppose the inequality holds for $\alpha$ and consider ${\widetilde{\alpha}}=\alpha\sqcup\{s\}$. Denote ${\widetilde{\beta}}_{2}=\beta_{2}\sqcup\{s\}$. Then
$\beta_{1}+({\widetilde{\alpha}}\sqcup\beta_{2})=\beta_{1}+(\alpha\sqcup{\widetilde{\beta}}_{2})\geq\alpha\sqcup(\beta_{1}+{\widetilde{\beta}}_{2})\geq\alpha\sqcup\{s\}\sqcup(\beta_{1}+\beta_{2})={\widetilde{\alpha}}\sqcup(\beta_{1}+\beta_{2}),$
where the first inequality follows by the inductive step and the second by the case of a singleton. This proves the case of three partitions.
The inequality for four partitions (4.4) follows by induction on the length of $\alpha_{1}$. If $\alpha_{1}=\emptyset$, this is the case of three partitions. Suppose $\alpha_{1}=(a_{1}\geq a_{2}\geq\dots\geq a_{k}>0)$ is a partition and let $\bar{\alpha}_{1}=(a_{1}\geq a_{2}\geq\dots\geq a_{k-1})$ so that $\alpha_{1}=\bar{\alpha}_{1}\sqcup\{a_{k}\}$. By induction
$(\bar{\alpha}_{1}\sqcup\beta_{1})+(\alpha_{2}\sqcup\beta_{2})\geq(\bar{\alpha}_{1}+\alpha_{2})\sqcup(\beta_{1}+\beta_{2}).$
Let us look at the contribution of $a_{k}$ on both sides. On the right hand side $a_{k}$ is added to the $k$-th part $a_{k}^{\prime}$ of $\alpha_{2}$ (written in decreasing order). In the left hand side, if we write $\alpha_{2}\sqcup\beta_{2}=(c_{1}\geq c_{2}\geq\dots)$, then $a_{k}$ is added to $c_{j}$ $j\geq k$, such that $c_{i}\geq a_{k}^{\prime}$ for all $i\leq j$. This means that the addition of $a_{k}$ in the two sides preserves the dominance order from before, and the claim follows inductively. ∎
@@ 4.3. Symplectic groups
At one extreme suppose that only endoscopic unitary factors appear, that is only polynomials $P$ with $\deg P>1$. Hence ${\mathbb{O}}^{\vee}_{\pi}$ is obtained by appending $2k_{P}$ copies of $m_{P}^{(1)}+m_{P}^{(2)}]$ and $2k_{P}$ copies of $|m_{P}^{(1)}-m_{P}^{(2)}|-1]$, for each $k_{P}>0$. On each $Sp(2n_{i})$, $i=1,2$, factor, we have
@@
$j_{\prod_{P}S_{a_{P}}^{k_{P}}}^{W_{n}}\left(\boxtimes_{P}(m_{P},\dots,2,1)^{\boxtimes k_{P}}\right).$
@@
$j_{\prod_{P}S_{k_{P}a_{P}}}^{W_{n}}\left(\boxtimes_{P}(m_{P}k_{P},\dots,2k_{P},k_{P})\right)=j_{\prod_{P}W_{k_{P}a_{P}}}^{W_{n}}(\sigma_{P}),$ (4.5)
@@
$\sigma_{P}=\begin{cases}(m_{P}k_{P},\dots,3k_{P},k_{P})\times((m_{P}-1)k_{P},\dots,4k_{P},2k_{P}),&m_{P}\text{ odd}\\ (m_{P}k_{P},\dots,4k_{P},2k_{P})\times((m_{P}-1)k_{P},\dots,3k_{P},k_{P}),&m_{P}\text{ even}.\end{cases}$
@@
$j_{\prod_{P}S_{a_{P}}^{k_{P}}}^{W_{n}}\left(\boxtimes_{P}(m_{P},\dots,2,1)^{\boxtimes k_{P}}\right)$
@@
By the calculation leading up to (4.6), we need to show that the nilpotent orbit for $j_{\prod_{P}W_{k_{P}a_{P}}}^{W_{n}}(\boxtimes_{P}\sigma_{P})$ is as claimed in the statement of the lemma. Since $j$-induction works in stages, this comes down to the case of two polynomials $P_{1},P_{2}$, that is to the easy verification that $j_{W_{1}\times W_{2}}^{W}(\sigma_{1}\boxtimes\sigma_{2})$ is the Springer representation for the nilpotent $(2m_{1}k_{1},\dots,4k_{1},2k_{1})+(2m_{2}k_{2},\dots,4k_{2},2k_{2})$. ∎
@@
    ${\mathbb{O}}^{\vee}_{\pi}=[1]\sqcup\bigsqcup_{P}[m_{P}^{(1)}+m_{P}^{(2)}]^{2k_{P}}\sqcup[|m_{P}^{(1)}-m_{P}^{(2)}|-1]^{2k_{P}}$.
@@
    $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\bigsqcup_{i=1,2}\left(\sum_{P}(2m_{P}^{(i)}k_{p},\dots,4k_{P},2k_{P})\right)$
@@
Claim (1) is just a restatement of the Lust–Stevens parametrization. Notice the presence of the entry $[1]$ because of the step (2s) in the procedure. Claim (2) is directly implied by Lemma 4.4. It remains to verify (3).
@@
${\mathbb{O}}^{\vee}_{\pi}=[1]\sqcup\bigsqcup_{P}(\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}}),$
@@
$^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\sum_{P}(2m_{P}k_{p},\dots,4k_{P},2k_{P})$
The combinatorial formula for computing $d^{\vee}$ on partitions $\underline{p}$ of $2n+1$ of odd orthogonal type is:
$d^{\vee}(\underline{p})=((\underline{p}^{-})_{C})^{t}=((\underline{p}^{t})^{-})_{C},$ (4.7)
where:
  • $\underline{p}\to\underline{p}^{-}$ decreases the smallest entry in $\underline{p}$ by $1$, thus making $\underline{p}^{-}$ a partition of $2n$;
  • $\underline{p}^{-}\to(\underline{p}^{-})_{C}$ is the $C$-collapse, resulting in a symplectic partition, see [CM, §6.3];
  • t is the transpose partition.
In our case, notice that if we denote the partition of $\mathbb{O}^{\vee}_{\pi}$ by $\underline{p}$, then $\underline{p}^{-}$ simply removes the entry $1$, and the $C$-collapse doesn’t do anything since $\underline{p}^{-}$ is already a symplectic partition.
Moreover
$(\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}})^{t}=(2m_{P}k_{p},\dots,4k_{P},2k_{P})$
which proves that $d^{\vee}({\mathbb{O}}_{\pi}^{\vee})={}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ if there is a single $P$. If there is more than one polynomial $P$, the claim follows as well by the compatibility (4.3) for two (or more) partitions $\alpha,\beta$.
@@
$d^{\vee}([1]\sqcup[m^{(1)}+m^{(2)}]^{2k}\sqcup[|m^{(1)}-m^{(2)}|-1]^{2k})$ $=\left((m^{(1)}+m^{(2)},m^{(1)}+m^{(2)}-2,\dots,m^{(1)}-m^{(2)}+2\right$.,
@@
$=(2m^{(1)}k,\dots,4k,2k)\sqcup(2m^{(2)}k,\dots,4k,2k),$
@@
$=\sum_{P}\left((2m_{P}^{(1)}k_{P},\dots,4k_{P},2k_{P})\sqcup(2m_{P}^{(2)}k_{P},\dots,4k_{P},2k_{P})\right)$
Using Lemma 4.3(5) and applying the inequality (4.4) repeatedly, one can see that $d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ is greater in the partition order than $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\bigsqcup_{i=1,2}\left(\sum_{P}(2m_{P}^{(i)}k_{p},\dots,4k_{P},2k_{P})\right)$. Notice that the difference between the two formulae is simply that $\Sigma$ and $\sqcup_{i=1,2}$ are interchanged.